OpenStudy (anonymous):

Please help. I'm confused. I need to simplify 3x - 5y = 15 into slope intercept form so that I can graph it. When I tried it myself I got y = 3/5x - 3, but when I looked it up to check my answers, a lot of people were saying that it's y = -3/5x - 3. Can you please solve it and go through the steps to explain it to me?

OpenStudy (anonymous):

3x - 5y = 15 -3x -3x -5y = -3x + 15 divide both sides by -5. y = (3/5)x -3 You are right

OpenStudy (anonymous):

It's likely they forgot that -3/(-5) = (3/5) since negatives cancel.

OpenStudy (anonymous):

Okay, thanks so much! You deserve a medal :)